A company has expected return of 15.6% and the market risk premium is 9%, with risk free rate being 3%. What is the beta for the company?
Answer #1

Expected return=risk free rate+beta*market risk premium

15.6=3+beta*9

Beta=(15.6-3)/9

=1.4
Answer #2

To calculate the beta for the company, we can use the Capital Asset Pricing Model (CAPM), which relates the expected return of an asset to its beta, the market risk premium, and the risk-free rate.

The formula for CAPM is:

Expected Return = Risk-Free Rate + Beta * Market Risk Premium

Given data: Expected Return (Company) = 15.6% Market Risk Premium = 9% Risk-Free Rate = 3%

Let's find the beta (β):

15.6% = 3% + β * 9%

Now, isolate β:

β * 9% = 15.6% - 3%

β * 9% = 12.6%

β = 12.6% / 9%

β ≈ 1.4

So, the beta for the company is approximately 1.4.


answered by: Hydra Master